# Learning phase transitions by siamese neural network

### v1

@@ 3.1 Test of the SNN framework
Figure 3: The output of SNN with the test anchor $p_{a}=0$ at $L=32$. The blue curve represents the similarity ($s$), while the ensemble average of the test samples is $1000$. The light blue curve represents the non-similarity ($1-s$).
According to the previous section, the entire SNN framework has been well trained by the set of samples within $p\in[0,0.1]\cup[0.9,1]$. To test the usability of SNN, firstly we select all samples of $p\in[0,0.1]$ as $S_{i}$ and only the configurations of anchor $p_{a}=0$ as $S_{j}$ at $L=32$ to form the input of test. The output $P_{0}$ is shown in the figure 3. For the similar curve $s$, $P_{0}$ is a stable value $1$ at $p\in[0,0.48]$ and 00 at $p\in[0.64,1]$, which means all configurations of anchor $p_{a}=0$ are $100\%$ having the same phase with samples of $p\in[0,0.48]$ but totally different with the samples of $p\in[0.64,1]$. In the area of $p\in(0.48,0.64)$, the similarity curve $s$ drops sharply, that is, the samples begin to be dissimilar to the ones of anchor $p_{a}=0$, and then it is completely in differ near $p=0.64$. The result of dissimilarity curve $(1-s)$ is opposite to the similarity. The evaluated critical point of SNN at $L=32$ is $p_{c}\simeq 0.6317$, which is close to the theoretical value $0.6447$ [35]. This shows that the SNN can relatively accurately predict the critical point of the (1+1) dimensional bond DP at $L=32$ by the selected anchor $p_{a}=0$
@@ 3.2 The dependence of anchor selection
Figure. 4 shows the similarity analysis of SNN with different anchors at $L=16,32,48$ and $64$. To study the dependence of anchors, we keep samples of $S_{i}$ from $p\in[0,1]$ but select anchor $p_{a}=0,0.4,0.8$ and $1$ for $S_{j}$ for comparison. Since during training process the SNN learns nothing about the samples of $p_{a}=0.4$ and $0.8$, here it also contains the idea of transfer learning, i.e., the range $p\in[0,0.1]\cup[0.9,1]$ as “source domain" and $p_{a}=0.4$ or $0.8$ as “target domain". As shown in Figure. 4, the similar curve $s$ changes with the different anchors. For $p_{a}=0$ (blue curve) and $0.4$ (brown curve), the similarity curve $s$ of all configurations is from “1" to “0" at $p\in[0,1]$, but for $p_{a}=0.8$ (red curve) and $1$ (green curve), it becomes the opposite trend as from “0" to “1". Because the configurations of $p_{a}=0$ and $0.4$ have similar phase to the ones of $p\in[0,0.1]$, but totally different at $p_{a}=0.8$ and $1$.
@@ 3.3 Extrapolated results of critical points
Figure 5: Extrapolation of the critical probability $p_{c}$ to infinite lattice size with anchor (a) $p_{a}=0$, (b) $p_{a}=0.4$, (c) $p_{a}=0.8$ and (d) $p_{a}=1$. The predictions of this $p_{c}^{\infty}$ are summarized in Table. 1.
To deeply analyze the dependence of the critical point predicted by SNN on anchor points, we extrapolate the results at $L=8,16,32,48$ and $64$ to zero on the $1/L$ scale under anchor points $p_{a}=0,0.4,0.8$ and $1$, by finite-size scale fitting (FSS) [48, 49, 50, 51]. As shown in Figure. 5, the critical values increase with the size of $L$, which is consistent with the layer of phase transition models under limited scale. The SNN results of $p_{c}$ at infinite lattice size are summarized in Table. 1. Comparing with the theoretical value $p_{c}=0.6447$ [35], it shows the SNN can relatively predict critical value $p_{c}^{\infty}$ within deviation. The closest result of SNN is $p_{c}=0.6426$ at anchor $p_{a}=0.4$, and the SNN results increase with the value of selected anchors, which also indicates the results of SNN may have anchor dependence. However, this achor dependence can be diluted by expanding the training set, which will be discussed in Subsection 4.2.
@@
for $1-s$ or $s$, where $\sigma$ is the scaling width. As shown in Figure. 6, by the technique of data collapse we can obtain the spatial correlation exponent $\nu_{\perp}$ via the scaling $(p-p_{c})L^{1/\nu_{\perp}}$. From $P_{0}$ at lattice sizes of $L=8,16,32,48$ and $64$, a proper value of $\nu_{\perp}$ can be found to fit the scaling. Table. 1 shows the fitted $\nu_{\perp}$ of each fixed anchors, all of which are consistent with the theoretical value $1.09$ [35]. Since $\nu_{\perp}$ has more dependence on raw configuration and also has lower precision, it is not as sensitive to changes with the anchors as $p_{c}^{\infty}$.
@@ 4 Discussion
4.1 The anchor point $p_{a}=0.64$
Figure 7: The output of SNN with anchor $p_{a}=0.64$ at $L=8,16,32,48$ and $64$.
[…]
@@ 4.2 The optimal training set
In previous tests, we sampled configurations $S_{i}$ and $S_{j}$ in the range of $p\in[0,0.1]\cup[0.9,1]$ as the training set of SNN, to minimize human intervention. However, from Section 3 it is obviously that the expected critical region is very narrow, which makes it possible to expand the training set. Typically, a wider and more accurate label training set is beneficial to improving the accuracy of NN predictions. Inspired by Ref. [20], an iterative approach is introduced to find an optimal interval of the training set. By defining the initial training set as $[0,l^{(0)}]\cup[r^{(0)},1]$ with $l^{(0)}=0.1$ and $r^{(0)}=0.9$, the updating process can be started with the initial estimate critical probability $p_{c}^{(0)}$:
@@
In Fig. 8, we show the iteration process to achieve the estimate critical value $p_{c}=0.6335$ of optimal training set $[0,0.58]\cup[0.69,1]$ for $L=32$, with anchor $p_{a}=0$. Obviously, the optimal training interval is much larger than the initial one, and the final result would also be more convinced. Table 2 gives the number of iterations, the optimal training interval and the critical probability predicted by SNN of $p_{a}=0$ for lattice size $L=8,16,32,48$ and $64$. Then, the critical probability of infinite size can also be obtained by the extrapolation in Fig. 9, as $p^{\infty}_{c}=0.6420$, and also by the data collapse we can get $\nu_{\perp}\simeq 1.09$. Table 3 shows the critical probabilities predicted by the optimal training interval under anchor point $p_{a}=0,0.4,0.8$ and $1$. Compared to the results of initial training interval in Table 1, a larger training set makes the SNN learn features more accurate, from which the results are less affected by the selected anchor values. Based on the previous summary, it can be inferred that by using the optimal training interval, selecting any anchor in training interval as test sample set can predict a reliable result. On the other hand, Table 3 also lists the results of supervised learning [47], autoencoder (AE) [47] and DANN [20] for comparison.
Figure 8: Evolution of the training set, and the corresponding critical bond probability values.
@@ 5 Conclusion
In this paper, we have applied the Siamese Neural Network (SNN) on the (1+1)-dimensional bond directed percolation to study its critical behaviors of nonequilibrium system. Here the SNN is used as a semi-supervised learning method, in which the key information of critical behavior can be predicted by labeling only a part of raw configurations as input for training. And the output of the SNN is similarity, which transforms the classification problems of phase transition into a discussion of difference, broadening the application of Machine learning (ML) techniques in the study of critical behaviors.
By the output similarity of SNN, we predict the critical points of the system at different lattice sizes, and then use the fitting based on finite-size scaling to to obtain the theoretical values at unlimited system. Further, the curve of similarity also can be fit to a sigmoid-like function, which allows us to calculate the spatial correlation exponent $\nu_{\perp}$of non-equilibrium phase transition system through data collapse.
@@ Acknowledgements
This work was supported in part by National Research Incubation Fund of Baoshan University(BYPY202216), the Baoshan University Doctoral Research Initiation Fund Project(BSKY202305), Academician workstation of Wenbangchun, Yunnan Province(Grant No.202205AF150032), Youth project of Yunnan Province Science and Technology Department(202401AU070035), Key Laboratory of Quark and Lepton Physics (MOE), Central China Normal University(Grant No.QLPL2022P01), and China Scholarship Council (CSC) - Swansea University scholarship.

### v2

@@ 4.1 Test of the SNN framework
Figure 4: The output of SNN with the test anchor $p_{a}=0$ at $L=32$. The blue curve represents the similarity ($s$), while the ensemble average of the test samples is $1000$. The light blue curve represents the non-similarity (1s1-s).
According to the previous section, the entire SNN framework has been well trained by the set of samples within $p\in[0,0.1]\cup[0.9,1]$. To test the usability of SNN, firstly we select all samples of $p\in[0,0.1]$ as $S_{i}$ and only the configurations of anchor $p_{a}=0$ as $S_{j}$ at $L=32$ to form the input of test. The output $P_{0}$ is shown in the Fig. 4. For the similar curve $s$, $P_{0}$ is a stable value $1$ at $p\in[0,0.48]$ and 0 at $p\in[0.64,1]$, which means all configurations of anchor $p_{a}=0$ are $100\%$ having the same phase with samples of $p\in[0,0.48]$ but totally different with the samples of $p\in[0.64,1]$. In the area of $p\in(0.48,0.64)$, the similarity curve $s$ drops sharply, that is, the samples begin to be dissimilar to the ones of anchor $p_{a}=0$, and then it is completely in differ near $p=0.64$. The result of dissimilarity curve (1s)(1-s) is opposite to the similarity. The evaluated critical point of SNN at $L=32$ is $p_{c}\simeq 0.6317$, which is close to the theoretical value $0.644700185(5)$ [54]. This shows that the SNN can relatively accurately predict the critical point of the (1+1) dimensional bond DP at $L=32$ by the selected anchor $p_{a}=0$ .
@@ 4.2 The dependence of anchor selection
Fig. 5 shows the similarity analysis of SNN with different anchors at $L=16,32,48$ and $64$. To study the dependence of anchors, we keep samples of $S_{i}$ from $p\in[0,1]$ but select anchor $p_{a}=0,0.4,0.8$ and $1$ for $S_{j}$ for comparison. Since during training process the SNN learns nothing about the samples of $p_{a}=0.4$ and $0.8$, here it also contains the idea of transfer learning, i.e., the range $p\in[0,0.1]\cup[0.9,1]$ as "source domain" and $p_{a}=0.4$ or $0.8$ as "target domain". As shown in Fig. 5, the similar curve $s$ changes with the different anchors. For $p_{a}=0$ (blue curve) and $0.4$ (brown curve), the similarity curve $s$ of all configurations is from "1" to "0" at $p\in[0,1]$, but for $p_{a}=0.8$ (red curve) and $1$ (green curve), it becomes the opposite trend as from "0" to "1". Because the configurations of $p_{a}=0$ and $0.4$ have similar phase to the ones of $p\in[0,0.1]$, but totally different at $p_{a}=0.8$ and $1$.
@@ 4.3 Extrapolated results of critical points
Figure 6: Extrapolation of the critical probability $p_{c}$ to infinite lattice size with anchor (a) $p_{a}=0$, (b) $p_{a}=0.4$, (c) $p_{a}=0.8$ and (d) $p_{a}=1$. The predictions of this $p_{c}^{\infty}$ are summarized in Table. 1. The results for each system size are obtained by running the SNN five times independently. The confidence interval for $p_{c}^{\infty}$ is determined by the Weighted Least-Squares Estimation method.
To deeply analyze the dependence of the critical point predicted by SNN on anchor points, we extrapolate the results at $L=8,16,32,48$ and $64$ to zero on the $1/L$ scale under anchor points $p_{a}=0,0.4,0.8$ and $1$, by finite-size scale fitting (FSS) [57, 58, 59, 60]. As shown in Fig. 6, the critical values increase with the size of $L$, which is consistent with the layer of phase transition models under limited scale. The SNN results of $p_{c}$ at infinite lattice size are summarized in Table. 1. Comparing with the theoretical value $p_{c}=0.644700185(5)$ [54], it shows the SNN can relatively predict critical value $p_{c}^{\infty}$ within deviation. The closest result of SNN is $p_{c}=0.6426$ at anchor $p_{a}=0.4$, and the SNN results increase with the value of selected anchors, which also indicates the results of SNN may have anchor dependence. However, this achor dependence can be diluted by expanding the training set, which will be discussed in Subsection 5.2.
@@
for 1s1-s or $s$, where $\sigma$ is the scaling width. As shown in Fig. 7, by the technique of data collapse we can obtain the spatial correlation exponent $\nu_{\perp}$ via the scaling $(p-p_{c})L^{1/\nu_{\perp}}$. From $P_{0}$ at lattice sizes of $L=8,16,32,48$ and $64$, a proper value of $\nu_{\perp}$ can be found to fit the scaling. Table. 1 shows the fitted $\nu_{\perp}$ of each fixed anchors, all of which are consistent with the theoretical value $1.09$ [39]. Since $\nu_{\perp}$ has more dependence on raw configuration and also has lower precision, it is not as sensitive to changes with the anchors as $p_{c}^{\infty}$.
@@
4.4 The SNN results of Ising model
(a) Different $p_{a}$
(b) $p_{a}=1.02$
(c) $p_{a}=1.02$
(d) $p_{a}=2.269$
Figure 8: The SNN results of Ising model. (a) The output of SNN with the test anchor $p_{a}=1.02$ when $L=32$, (b) Extrapolation of the critical probability $T_{c}$ to infinite lattice size with anchor (a) $p_{a}=1.02$, (c) The collapse of the average output layer as a function of $(T-T_{c})L^{1/\nu_{\perp}}$, with anchor $p_{a}=1.02$, (d) The output of SNN with anchor $p_{a}=2.269$ at $L=8,16,32,48$ and $64$.
To evaluate the cross-system robustness of the SNN method, we apply it to the two-dimensional Ising model. Fig. 8 presents the results of the Ising model analysis using the SNN. Similar to the SNN learning approach for DP, we choose different anchor points of temperature. To simplify the analysis, we select anchor points as $p_{a}=1.02,1.52,3.52$ and $4.02$, which are the boundary values of the entire labeled temperature range $T\in[1.02,1.52]\cup[3.52,4.02]$.
In Fig. 8 (a) shows the results of SNN for the four chosen anchor points at $L=32$. Obviously, according to the studies in the previous subsections, these outputs are in line with expectations: the similarity $s$ increases to $1$ as the two compared configurations are from closer temperature values and it decreases to 0 on the contrary. The black dash line in Fig. 8(a) represents the theoretical critical value of the Ising model at $2.269$ [42]. Due to finite-size effects, it is a discrepancy between the intersection of the curves and the theoretical value.
To obtain the predicted critical value $T_{c}$ from the SNN, we perform the finite-size scaling analysis for the anchor point at 1.02 with system sizes of $L=8,16,32,48$, and $64$, as we have used for the (1+1) DP before. The results are shown in Fig. 8(b), and the predicted critical temperature of infinite system is $T_{c}^{\infty}=2.2777\pm 0.0033$, close to the theoretical one. And also, by the data collapse method, we achieve a good fitting of the similarity curves for the five system sizes as illustrated in Fig. 8(c), and then obtain the correlation exponent $\nu=1.0$ as the numerical solution in Ref. [61]. In Fig. 8(d), we present the SNN results by choosing a specific anchor point $p_{a}=2.269$, close to the theoretical critical temperature. The similarity output exhibits a peak at the anchor point, indicating that at this temperature the input configurations are only similar to themselves. It is an interesting behavior and we will give detailed discussions in Section 5.1.
@@ 5 Discussion
5.1 In-depth study on the selection of anchor points
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 9: The SNN results of the (1+1) dimensional directed percolation model by choosing different anchor points at $L=32$. (a) $p_{a}<p_{c}$, (b) $p_{a}$ approaches $p_{c}$ , (c) $p_{a}>p_{c}$. The black dashed line represents the theoretical critical value $p_{c}=0.644700185(5)$.
To deeply discover the relationship between the predicted critical point and the choice of anchor points, here we conducted more tests using different anchor points, taking the (1+1) dimensional directed percolation model as an example. In section 4.2, we have selected four points deviating from the critical point as anchor points $p_{a}=0,0.4,0.8,$ and $1$ to determine the system’s critical value and spatial correlation exponent through finite-size scaling and data collapse. Now to observe the impact of anchor points on critical value prediction over a broader range, we supplement with additional anchor points $p_{a}=0.1,0.2,0.3,0.4,$ and $0.5$, as shown in Fig. 9. For the results of SNN, it can be seen that the similarity $s$ decreases progressively from $p=0$ to $1$, and the outputs of five curves are nearly consistent. From Fig.  9, it is evident that the critical points predicted by the SNN closely match the theoretical value $p_{c}=0.644700185(5)$ [54]. As in Section  4.2, by the skill of finite-size scaling and data collapse, we also can obtain the critical value and spatial correlation exponent of infinite system. In Fig. 9(c), by using $p_{a}=0.75,0.8,0.85,0.9$ and 0.95, the SNN also can gives the critical information of system, similar to Fig. 9(a).
In contrast, Fig. 9(b) displays entirely different behavior compared to Fig. 9(a) and 9(c). In Fig. 9(b), we chose anchor points at $p_{a}=0.62,0.63,0.64,0.65,$ and $0.66$, in the critical region. In these cases, the curves $s$ are not sigmoid-like function but have a structure of single peak. The maximum values of the peak for each lattice size $L$ are all close to 1. The reason of this behavior is that the configurations close to the critical point are only very similar to themselves but different from the ones of other bond probabilities. As the value of $p$ moves away from the anchor $p_{a}=0.62,0.63,0.64,0.65$ and $0.66$, the similarity curves decrease rapidly within a short distance in the range of critical area and then changes slowly. Because of fluctuations near the critical region in the system, the SNN results also show a degree of deviation from the perspective of $x$-coordinates pp.
The above analysis demonstrates that it is very challenging to obtain critical information of the system by using points in the critical region as anchor points. Hence, in the main text, we opted for points far from the critical region as anchor points.
@@ 5.2 The optimal training set
In previous tests of the (1+1) dimensional directed percolation model, we sampled configurations $S_{i}$ and $S_{j}$ in the range of $p\in[0,0.1]\cup[0.9,1]$ as the training set of SNN, to minimize human intervention. However, from Section 4 it is obviously that the expected critical region is very narrow, which makes it possible to expand the training set. Typically, a wider and more accurate label training set is beneficial to improving the accuracy of NN predictions. Inspired by Ref. [22], an iterative approach is introduced to find an optimal interval of the training set. By defining the initial training set as $0,l^{(0)}]\cup[r^{(0)},1]$ with $l^{(0)}=0.1$ and $r^{(0)}=0.9$, the updating process can be started with the initial estimate critical probability $p_{c}^{(0)}$:
@@
In Fig. 10, we show the iteration process to achieve the estimate critical value $p_{c}=0.6335$ of optimal training set $0,0.58]\cup[0.69,1]$ for $L=32$, with anchor $p_{a}=0$. Obviously, the optimal training interval is much larger than the initial one, and the final result would also be more convinced. Table 2 gives the number of iterations, the optimal training interval and the critical probability predicted by SNN of $p_{a}=0$ for lattice size $L=8,16,32,48$ and $64$. Then, the critical probability of infinite size can also be obtained by the extrapolation in Fig. 11, as $p^{\infty}_{c}=0.6420$, and also by the data collapse we can get $\nu_{\perp}\simeq 1.09$ in Fig. 11. Table 3 shows the critical probabilities predicted by the optimal training interval under anchor point $p_{a}=0,0.4,0.8$ and $1$. Compared to the results of initial training interval in Table 1, a larger training set makes the SNN learn features more accurate, from which the results are less affected by the selected anchor values. Based on the previous summary, it can be inferred that by using the optimal training interval, selecting any anchor in training interval as test sample set can predict a reliable result. On the other hand, Table 3 also lists the results of supervised learning [53], autoencoder (AE) [53] and DANN [22] for comparison.
Figure 10: Evolution of the training set, and the corresponding critical bond probability values with DP system size $L=32$
@@ 6 Conclusion
In this paper, we have applied a Siamese Neural Network (SNN) to investigate the critical behavior of two prominent phase transition models: the (1+1) dimensional bond directed percolation of non-equilibrium system, and the 2-dimensional Ising model as a classic example of equilibrium phase transition. Here the SNN is used as a semi-supervised learning method, in which the key information of critical behavior can be predicted by labeling only a part of raw configurations as input for training. And the output of the SNN is similarity, which transforms the classification problems of phase transition into a discussion of difference, broadening the application of Machine learning (ML) techniques in the study of critical behaviors.
By the output similarity of SNN, we predict the critical points of the system at different lattice sizes, and then use the fitting based on finite-size scaling to to obtain the theoretical values at unlimited system. Further, the curve of similarity also can be fit to a sigmoid-like function, which allows us to calculate the spatial correlation exponent $\nu_{\perp}$ of the (1+1) dimensional bond directed percolation and correlation exponent $\nu$ of the 2- dimensional Ising model in phase transition systems through data collapse.
@@ Acknowledgements
This work was supported in part by Yunnan Fundamental Research Projects (Grant 202401AU070035), the Baoshan University Doctoral Research Initiation Fund Project(BSKY202305), and China Scholarship Council (CSC) - Swansea University scholarship.